OpenStudy (anonymous):

What are the exact solutions of x2 − 5x − 7 = 0? x = the quantity of negative 5 plus or minus the square root of 3 all over 2 x = the quantity of 5 plus or minus the square root of 3 all over 2 x = the quantity of negative 5 plus or minus the square root of 53 all over 2 x = the quantity of 5 plus or minus the square root of 53 all over 2

OpenStudy (debbieg):

What do you get when you factor that trinomial?

OpenStudy (anonymous):

idk

OpenStudy (skullpatrol):

Have you seen the "quadratic formula?"

OpenStudy (debbieg):

It's just "un-FOILing". You need to write \(x^2 − 5x − 7 \) as a product (multiplication) of two binomials: \(\Large x^2 − 5x − 7 =(....?....)(....?....)\) so you need to figure out what goes in the ( )'s. Since the trinomial starts with an \(x^2\) without a coefficient in front of it, you KNOW it has to be: \(\Large x^2 − 5x − 7 =(x\pm ?)(x\pm ?)\) So all you need is to figure out the LAST terms, and the signs in the middle.

OpenStudy (debbieg):

oh wait, may bad.... @skullpatrol is right. Need the QF on this one.

OpenStudy (debbieg):

*my

OpenStudy (debbieg):

I hadn't looked at the answers... lol.

OpenStudy (anonymous):

still don't get it

OpenStudy (debbieg):

You have \(\Large x^2 − 5x − 7 =0\) That's of the form \(\Large ax^2 +bx +c =0\) For YOUR equation, what is a, b and c? a= ? b= ? c= ?

OpenStudy (anonymous):

okay

OpenStudy (skullpatrol):